Exercise 12-12 Volume Trade-Off Decisions [LO12-5]

Benoit Company produces three products—A, B, and C. Data concerning the three products follow (per unit):

Product

A B C
Selling price $

90.00

$ 60.00 $ 80.00
Variable expenses:
Direct materials 27.00 18.00 9.00
Other variable expenses 27.00 27.00 47.00
Total variable expenses 54.00 45.00 56.00
Contribution margin $ 36.00 $ 15.00 $ 24.00
Contribution margin ratio 40 % 25 % 30 %

The company estimates that it can sell 750 units of each product per month. The same raw material is used in each product. The material costs $3 per pound with a maximum of 5,400 pounds available each month.

Required:

1. Calculate the contribution margin per pound of the constraining resource for each product.

2. Which orders would you advise the company to accept first, those for A, B, or C? Which orders second? Third?

3. What is the maximum contribution margin that the company can earn per month if it makes optimal use of its 5,400 pounds of materials?
Answer #1

1) Contribution margin per pound:

A

B

C

Contribution margin per unit   $

36

15

24

Pound per unit A=27/3,B=18/3,C=9/3

9

6

3

Contribution margin per pound

4

2.5

8

Ranking

2

3

1

2) Optimum use

Product

rank

Product A

Second

product B

Third

Product C

First

3) Maximum contribution margin

Product

Pound available

Pounds consumed

Units produced

C

5400

750*3= 2250

750

A

5400-2250= 3150

3150

3150/9 pound =350

B

3600-3600=0

0

0

Contribution margin: [750*$36]+[0*$15]+[350*24]

= 27000+0+8400

= $ 27000